### v5

@@ 5.1. Healthcare
GenAI is transforming healthcare across clinical, biomedical, and administrative domains, but each application underscores the importance of RAI safeguards. For example, while generative models improve efficiency in clinical note generation and triage chatbots (Johnson and Patel, 2023; Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support. Disease detection and drug discovery pipelines (Miotto et al., 2022; Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment. Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases. Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024). Even promising breakthroughs, such as GenAI-designed therapeutics (Smith et al., 2025), requires human oversight on trial integrity, equitable access, and long-term monitoring. In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
RAI Challenges.
@@
Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019; McGraw and Mandl, 2020; Goldstein and Holtzman, 2021), hallucination of critical clinical facts (Menny and Braun, 2019; Mamo et al., 2023a; Li et al., 2023c), algorithmic bias in diagnostics (Obermeyer et al., 2019; Gonzalez et al., 2022; Venkatasubramanian, 2020), unclear accountability for errors (Kaminski, 2022; Kraft and Jonas, 2021), opacity of generative models (Ghorbani et al., 2019; Tjoa and Guan, 2023; Donoso-Guzmán et al., 2025), and difficulty of compliance with traditional regulations (Price and Gerke, 2022; Kaissis et al., 2021; Khalid et al., 2023). Empirical studies highlight these risks: a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023c); Rosbach et al. (2024) found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
RAI Strategies.
@@
Several strategies are being employed to enable responsible GenAI adoption in healthcare. Grounding methods such as retrieval-augmented generation (RAG) help reduce hallucinations by linking model outputs to verifiable medical sources (Wu et al., 2024). Human-in-the-loop (HITL) validation integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias (Stojanovic et al., 2021; He et al., 2023b). Privacy-preserving approaches, including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration (Kaissis et al., 2021; Khalid et al., 2023). Fairness-aware data augmentation and structured fairness audits are being applied to improve equity in clinical outcomes (Pfohl et al., 2022; Chen et al., 2023). Post-hoc explainability techniques that are applied after model training to make opaque AI models interpretable such as SHAP and LIME, are critical for clinician trust (Nasarian et al., 2023; Donoso-Guzmán et al., 2025; Guidotti et al., 2021). Alignment with regulatory frameworks like Food and Drug Administration (FDA) guidance ensures safe and accountable deployment (Kaminski, 2022; Kraft and Jonas, 2021).
@@ 5.2. Finance
GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Ngai et al., 2011; Cao, 2022b). Key applications include customer-facing bots that provide interactive banking and wealth-management support (Wang et al., 2023b; Chen and Lee, 2022; Ngai et al., 2011), modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Hand, 2018; Mehrabi et al., 2021b; Cao, 2022b), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (Zhang et al., 2023; Gupta et al., 2024; HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Silva et al., 2024; Mehrabi et al., 2021b; Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Araci, 2019; Li et al., 2023a; Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Patki et al., 2016; Wiese et al., 2020; Cao, 2022b). Empirical deployments highlight these advances. For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011). Zest AI raised loan approvals for minority applicants by 15% in a 2024 U.S. bank case study (Hand, 2018; Mehrabi et al., 2021b). Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b). JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023). HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025; Cao, 2022b). Silva et al. (2024) reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
RAI Challenges.
@@
GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption. These include discriminatory lending that replicates historical bias in credit and valuation data (Hand, 2018; Mehrabi et al., 2021b; Adams et al., 2022), security vulnerabilities where adversarial manipulation targets fraud detection or trading systems (Biggio et al., 2013; Kraft and Jonas, 2021; Ngai et al., 2011), consumer data privacy concerns arising from reliance on sensitive financial records (McGraw and Mandl, 2020; Khalid et al., 2023; Goldstein and Holtzman, 2021), opacity of GenAI that limits auditability and accountability (Doshi-Velez and Kim, 2017; Guidotti et al., 2021; Office of the Comptroller of the Currency, 2020), hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice (Li et al., 2023c; Mamo et al., 2023a), and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (U.S. Securities and Exchange Commission, 2021; Hendry et al., 2022). Empirical use cases underscore these risks. For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022). Citibank was fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020). The Robinhood–GameStop episode demonstrated how synthetic order-flow simulations amplified volatility and required regulatory intervention (U.S. Securities and Exchange Commission, 2021).
RAI Strategies.
@@
Multiple strategies are being prioritized to address these financial risks. Structured fairness audits and fairness-aware data augmentation are increasingly required in lending and credit scoring (Pfohl et al., 2022; Chen et al., 2023; Mehrabi et al., 2021b). Post-hoc explainability methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models (Guidotti et al., 2021; Nasarian et al., 2023; Donoso-Guzmán et al., 2025). Privacy-preserving techniques, including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data (Kaissis et al., 2021; Khalid et al., 2023; McGraw and Mandl, 2020). Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets (Patki et al., 2016; Wiese et al., 2020). Alignment with regulatory compliance frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022; Kraft and Jonas, 2021; Office of the Comptroller of the Currency, 2020). Finally, embedding HITL oversight across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions (Stojanovic et al., 2021; He et al., 2023b; Silva et al., 2024).
@@ 5.3. Education
GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Luckin et al., 2022; Holmes et al., 2022). Key applications include automated grading that accelerates evaluation and reduces instructor workload (Zawacki-Richter et al., 2019; Holmes et al., 2022), personalized tutoring systems that adapt to student knowledge gaps and learning styles (Kulik and Fletcher, 2023; Carnegie Learning, 2023), content generation that produces adaptive curricula and study materials (Kasneci et al., 2023; Zhang and Zhao, 2024), accessibility tools that deliver multimodal and multilingual support for inclusive education (Almalki and Khan, 2023; Holstein and colleagues, 2023), and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024). Recent deployments provide evidence of these advances. Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S. high schools in 2023 (Carnegie Learning, 2023). A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024). A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024). A U.K. school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
RAI Challenges.
@@
Responsible use of GenAI in education faces several pressing concerns. These include cultural insensitivity where models fail to capture local norms and values (Kasneci et al., 2023; UNESCO, 2021), student data privacy and particularly minors’ data privacy given the sensitivity of educational records (Holmes et al., 2022; Livingstone et al., 2021), skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving (Luckin et al., 2022; Lai and Xu, 2024), and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Holstein and colleagues, 2023; Zawacki-Richter et al., 2019). Empirical use cases underscore these risks. For example, India’s government-run “Newton” homework bot trained only on U.S. and U.K. corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022). A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023). In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
RAI Strategies.
@@
Multiple strategies are being explored to ensure responsible generative AI in education. Fairness audits and external ethics audits verify alignment with frameworks such as the UNESCO AI-ethics rubric and can keep bias metrics below accepted thresholds (UNESCO, 2021; Pfohl et al., 2022; Chen et al., 2023). Privacy-preserving techniques including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021; Khalid et al., 2023). Clear regulation and policy guidelines developed by international and national bodies provide governance and accountability mechanisms (UNESCO, 2021; Holmes et al., 2022). Finally, academic integrity tools such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Kasneci et al., 2023; Zhang and Zhao, 2024).
@@ 5.4. Arts and Entertainment
GenAI is reshaping creative production and cultural participation in the arts and entertainment sector (Crawford, 2021; Traitter and Klein, 2022). Prominent applications include creative writing tools that assist in script, story, and lyric development (McStay, 2023; Zhang and Lee, 2023), generative art and design platforms that expand aesthetic experimentation (Elgammal et al., 2019; Schneider and Jörgens, 2022), music composition and production systems that automate melody, arrangement, and mastering (Davis et al., 2024; Briot et al., 2020), film production support for visual effects, dubbing, and storyboarding (Miranda and Perez, 2022; Rodriguez and Santos, 2024), personalized recommendation engines that curate media consumption at scale (Celma, 2010; Anderson and Lee, 2023), and virtual influencers and avatars that shape marketing and fan engagement (Traitter and Klein, 2022; Crawford, 2021). Recent deployments provide evidence of these advances. For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023). A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022). Davis et al. (2024) released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
RAI Challenges.
@@
Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns. These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization (U.S. Copyright Office, 2023; Samuelson, 2023), job displacement that threatens creative labor markets (Crawford, 2021; McStay, 2023), cultural insensitivity when generative tools fail to respect community values or norms (UNESCO, 2021; Kasneci et al., 2023), deepfakes and misinformation that erode authenticity and trust (Chesney and Citron, 2019; Fazio, 2023), authorship and agency disputes where creative ownership is contested (U.S. Copyright Office, 2023; Samuelson, 2023), insufficient consent in data usage for training models (WeTransfer, 2025; Crawford, 2021), and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019). Empirical use cases underscore these risks. For example, the SAG-AFTRA and WGA strikes of 2023 protested generative AI’s role in replacing talent and using likenesses without consent or compensation. The U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S. Copyright Office, 2023). The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries. In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
RAI Strategies.
@@
Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment. Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value (McStay, 2023; Crawford, 2021). Structured bias audits and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping (UNESCO, 2021; Pfohl et al., 2022). Deepfake detection tools are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019; Fazio, 2023). Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023). International ethical guidelines and transparency labeling requirements provide standards for disclosure and attribution (UNESCO, 2021; Samuelson, 2023). Finally, intellectual property management systems are being advanced to secure consent and compliance. For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these system requires explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.

### v6

@@ 5.1. Healthcare
Generative models improve efficiency in clinical note generation and triage chatbots (Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support. Disease detection and drug discovery pipelines (Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment. Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases. Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024). Even promising breakthroughs, such as GenAI-designed therapeutics, require human oversight on trial integrity, equitable access, and long-term monitoring. In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
ResGenAI challenges
@@
Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019), hallucination of critical clinical facts (Li et al., 2023b), algorithmic bias in diagnostics (Obermeyer et al., 2019), unclear accountability for errors (Kaminski, 2022), opacity of generative models (Ghorbani et al., 2019), and difficulty of compliance with traditional regulations (Price and Gerke, 2022). Empirical studies highlight these risks, for example, a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023b); Rosbach et al. further found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
Evaluation context
Mapping these risks to our rubric, the dominant healthcare concerns correspond to C4 (hallucination of clinical facts), C1 (diagnostic bias), and C6 (patient data leakage). Only 7 of 21 benchmarks in Table 4 score above zero on C4 (Misinformation/ Disinformation), and none specifically target medical hallucination. For privacy, only PrivLM-Bench and PrivacyLens provide strong C6 (Privacy) coverage, however, neither addresses HIPAA-specific leakage vectors such as clinical note memorization. C7 (System-level failure) is critical for RAG-based clinical pipelines and is tested by only 4 of 21 benchmarks. The KPIs most relevant to this domain are the high-stakes error rate $E_{h}$ (KPI 9), operationalized as hallucinated clinical facts per $N$ consultations, data privacy compliance $P$ (KPI 2) under HIPAA, and fairness metrics SPD/DI (KPIs 3–4) stratified by patient demographics.
ResGenAI strategies
@@
Several strategies are being employed to enable responsible GenAI adoption in healthcare. Grounding methods (C4, KPI 9) such as RAG can help reduce hallucinations by linking model outputs to verifiable medical sources. Effectiveness should be tracked via $E_{h}$; a reported 1.47% hallucination rate provides a baseline. Human-in-the-loop (HITL) validation (C7, C8, KPI 10) integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias. Audit frequency $F_{a}$ and resolution time $T_{r}$ should be reported.Privacy-preserving approaches (C6, KPI 2), including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration. Fairness-aware data augmentation and structured fairness audits (C1, KPIs 3–5) are being applied to improve equity in clinical outcomes. HumaniBench (C1=2 in Table 4) offers a starting template but lacks clinical subgroup stratification. Post-hoc explainability (C9, KPIs 7–8) techniques such as SHAP and LIME are critical for clinician trust.Explanation quality should be assessed via the local faithfulness gap LFG (KPI 8). Alignment with regulatory frameworks (C10) like FDA guidance ensures safe and accountable deployment. Table 6 shows that only HELM and MLCommons v0.5 provide partial transparency support relevant to FDA conformity assessment.
@@ 5.2. Finance
GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Cao, 2022). Key applications include customer-facing bots that provide interactive banking and wealth-management support, modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Cao, 2022), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Cao, 2022). Empirical deployments highlight these advances. For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011). Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b). JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023). HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025). Silva et al. also reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
ResGenAI challenges
@@
GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption. These include discriminatory lending that replicates historical bias in credit and valuation data , security vulnerabilities where adversarial manipulation targets fraud detection or trading systems, consumer data privacy concerns arising from reliance on sensitive financial records, opacity of GenAI that limits auditability and accountability, hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice, and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (Hendry et al., 2022). Empirical use cases underscore these risks. For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022). Citibank was also fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
Evaluation context
The primary finance risks map to C1 (lending discrimination), C3 (adversarial manipulation of trading/fraud systems), C6 (consumer data privacy), and C4 (hallucinated financial advice). No benchmark in Table 4 specifically targets lending discrimination or financial adversarial attacks. The most relevant KPIs are SPD and DI (KPIs 3–4) applied to loan approval rates across demographic groups, robustness $R$ (KPI 6) under financial adversarial scenarios, and audit frequency $F_{a}$ (KPI 10) aligned with SEC and OCC reporting cadences.
ResGenAI strategies
@@
Multiple strategies are being prioritized to address these financial risks. Structured fairness audits (C1, KPIs 3–4) and fairness-aware data augmentation are increasingly required in lending and credit scoring. Post-hoc explainability (C9, KPIs 7–8) methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models. Privacy-preserving techniques (C6, KPI 2), including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data. Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets. Alignment with regulatory compliance (C10) frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022). Finally, embedding HITL oversight (C7, KPI 10) across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions.
@@ 5.3. Education
GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Holmes et al., 2022). Key applications include automated grading that accelerates evaluation and reduces instructor workload, personalized tutoring systems that adapt to student knowledge gaps and learning styles, content generation that produces adaptive curricula and study materials, accessibility tools that deliver multimodal and multilingual support for inclusive education, and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024). Recent deployments provide evidence of these advances. Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S. high schools in 2023 (Carnegie Learning, 2023). A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024). A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024). A U.K. school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
ResGenAI challenges
@@
Responsible use of GenAI in education faces several pressing concerns. These include cultural insensitivity where models fail to capture local norms and values, student data privacy and particularly minors’ data privacy given the sensitivity of educational records , skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving, and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Zawacki-Richter et al., 2019). Empirical use cases underscore these risks. For example, India’s government-run “Newton” homework bot trained only on U.S. and U.K. corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022). A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023). In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
Evaluation context
Education risks concentrate on C6 (minors’ data privacy), C4 (hallucinated pedagogical content), C1 (cultural bias), and C7 (system-level failure in tutoring pipelines). No benchmark addresses C6 in the context of minors’ data, and the user inclusivity index $U$ (KPI 11) critical for accessibility in educational settings, lacks any standardized evaluation protocol. Skills-evasion effects (a form of C7 system-level failure) are entirely absent from current safety suites. Priority KPIs include $P$ (KPI 2) under COPPA/FERPA, $E_{h}$ (KPI 9) for pedagogically harmful outputs, and $U$ (KPI 11) for WCAG conformance.
ResGenAI strategies
@@
Multiple strategies are being explored to ensure responsible generative AI in education. Fairness audits (C1, KPIs 3–4) and external ethics audits verify alignment with frameworks such as the UNESCO(UNESCO, 2021a) AI-ethics rubric and can keep bias metrics below accepted thresholds. Privacy-preserving techniques (C6, KPI 2) including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021). Clear regulation and policy guidelines (C10) developed by international and national bodies provide governance and accountability mechanisms. Finally, academic integrity tools (C5, C9) such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Zhang and Zhao, 2024).
@@ 5.4. Arts and Entertainment
GenAI is reshaping creative production and cultural participation in the arts and entertainment sector. Prominent applications include creative writing tools that assist in script, story, and lyric development, generative art and design platforms that expand aesthetic experimentation, music composition and production systems that automate melody, arrangement, and mastering, film production support for visual effects, dubbing, and storyboarding, personalized recommendation engines that curate media consumption at scale and virtual influencers and avatars that shape marketing and fan engagement (Crawford, 2021). Recent deployments provide evidence of these advances. For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023). A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022). Davis et al. released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
ResGenAI challenges
@@
Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns. These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization, job displacement that threatens creative labor markets, cultural insensitivity when generative tools fail to respect community values or norms deepfakes and misinformation that erode authenticity and trust, authorship and agency disputes where creative ownership is contested, insufficient consent in data usage for training models, and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019). Empirical use cases underscore these risks. For example, the U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S. Copyright Office, 2023). The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries. In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
Evaluation context
The dominant risks map to C5 (deepfakes and media integrity), C1 (representational bias in generated content), and C9 (attribution transparency). Only SHIELD and MM-SafetyBench provides explicit C5 evaluation in Table 4, and its scope does not extend to voice cloning or music mimicry, the vectors most salient for this sector. Copyright memorization, a C6-adjacent concern, is not addressed by any privacy benchmark. Priority KPIs are currently underspecified for creative domains; provenance-verification rates and consent-compliance metrics (analogous to KPI 2) are needed but lack standardized protocols.
ResGenAI strategies
@@
Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment. Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value. Structured bias audits (C1, KPIs 3–4) and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping . Deepfake detection tools (C5) are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019). Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023). International ethical guidelines (C10) and transparency labeling requirements provide standards for disclosure and attribution. Finally, intellectual property management systems (C6, C9) are being advanced to secure consent and compliance. For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these systems require explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.